Please help..
Find an equation of the line having the given slope and containing the given point. m=7, (9,3)

Found 2 solutions by nerdybill, solver91311:
Answer by nerdybill(7384)

plug given data:
m=7, (9,3)
.
into "point-slope" form:
y - y1 = m(x - x1)
y - 3 = 7(x - 9)
y - 3 = 7x - 63
y = 7x - 60 (this is what they're looking for)


Answer by solver91311(24713)



Use the point-slope form of an equation of a line.



where are the coordinates of the given point and is the given slope.


John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism